OpenStudy (anonymous):

Given the following quadratic equation, determine if it has a maximum or a minimum value. Then find the maximum or minimum value. Please show all of your work. f(x) = = -x^2-4x-5

OpenStudy (anonymous):

maximum of -5 I used a TI-83 Plus I don't know how to manually solve this kind of st00f.

OpenStudy (anonymous):

Wait, it's actually a maximum of -2

OpenStudy (anonymous):

it has a max because the leading coefficient is negative ( it is -1)

OpenStudy (anonymous):

max is found by using \[x=-\frac{b}{2a}\] in this case it is \[x=-\frac{-4}{2\times -1}=-2\] replace x by -2 and you get \[-(-2)^2-4\times -2-5=-4+8-5=-1\]

OpenStudy (anonymous):

Thank you.